# Cascade-Zero123: One Image to Highly Consistent 3D with Self-Prompted Nearby Views

### v1

@@ Abstract
Synthesizing multi-view 3D from one single image is a significant and challenging task. For this goal, Zero-1-to-3 methods aim to extend a 2D latent diffusion model to the 3D scope. These approaches generate the target-view image with a single-view source image and the camera pose as condition information. However, the one-to-one manner adopted in Zero-1-to-3 incurs challenges for building geometric and visual consistency across views, especially for complex objects. We propose a cascade generation framework constructed with two Zero-1-to-3 models, named Cascade-Zero123, to tackle this issue, which progressively extracts 3D information from the source image. Specifically, a self-prompting mechanism is designed to generate several nearby views at first. These views are then fed into the second-stage model along with the source image as generation conditions. With self-prompted multiple views as the supplementary information, our Cascade-Zero123 generates more highly consistent novel-view images than Zero-1-to-3. The promotion is significant for various complex and challenging scenes, involving insects, humans, transparent objects, and stacked multiple objects etc.
Figure 1: Rather than adopting the one-to-one generation pipeline as in Zero-1-to-3 [29], Cascade-Zero123 progressively extracts the 3D information from one single image via self-prompted nearby views. View-consistent images can be generated by constructing the structure in a cascade manner. Cascade-Zero123 shows strong capability on various complex objects, e.g. insects, robots, or multiple objects stacked.
@@ 1 Introduction
Large latent diffusion models have achieved remarkable performance in synthesizing high-quality 2D images with significant zero-shot capability [46, 37, 48, 44]. To migrate such generalizability to the 3D field, many researchers have attempted to adapt diffusion models to 3D tasks [49, 18, 35, 23, 36, 57, 8, 55, 42, 26, 58, 5, 59, 65, 54, 71, 1, 20, 38, 14, 13, 33, 30, 69, 62, 12]. Generating the 3D object from a single image has become an emerging research topic in 3D computer vision, for its flexibility and convenience of various real-world applications [40, 45] including virtual reality, 3D content creation, computer gaming, movie or animation rendering, etc.
Regarding the task of using a single image to generate 3D, there are two main research directions. The first is to directly train a 3D model [4, 63, 22, 38, 21], but due to the significant resources required for collecting and organizing 3D data, large-scale 3D datasets are limited. As a result, methods that directly train 3D models perform well only on scenes similar to the training set and lack generalization ability. The second direction focuses on effectively utilizing the 2D latent diffusion, the highlight is to leverage the zero-shot generalization capability obtained from pre-training on large-scale 2D datasets. Models such as Zero-1-to-3 [29] are view-conditioned large latent diffusion models, in which the pre-training of multi-view images based on angle conditioning enables the models to have the ability to reconstruct the structure of unseen samples. For each input image, these models can directly output the corresponding novel-view images. However, the prior 2D diffusion models and view-conditioned latent diffusion models usually generate each image separately. As a result, the generated images are hard to be geometrically and visually consistent across different views. How to better utilize the capabilities of view-conditioned latent diffusion models and improve these pre-trained models to perform better novel view synthesis becomes a worthy explored direction.
In practice, for objects with simple structures, e.g. centrally symmetric or left-right symmetric objects, Zero-1-to-3 demonstrates excellent 3D cognitive ability and generates consistent novel views. However, for objects with complex structures, it is challenging to synthesize high-quality novel-view images when the camera pose changes dramatically and incurs completely unseen parts.
Considering the above issues, we propose a framework, named Cascade-Zero123, to reduce the difficulty of the task, which allows Zero-1-to-3 to progressively extract 3D information from a single image based on the self-prompted nearby views. Cascade-Zero123 consists of two cascade Zero-1-to-3 models called Base-0123 and Refiner-0123. First, we use Base-0123 to generate a set of nearby views, allowing the model to develop a basic 3D prior knowledge of the object. These self-prompt views are then used as condition images and inputted to Refiner-0123. Our pipeline decomposes the task of directly rotating an image to a large rotation angle. Instead, it transforms the task into two parts, which are generating multiple nearby images from the input image and utilizing them as prompts to generate the final target views. As shown in Fig. 2, due to the context cross-attention between multi-view self-prompted condition images and the noise latent in the second Zero-1-to-3, the model’s output of multi-view images gains a certain level of geometric and visual consistency.
@@
  • We propose a cascade framework to progressively extract 3D information from a single image, preventing directly rotating to completely unseen parts of objects. This can better excavate the novel view synthesis capability of Zero-1-to-3.
  • We introduce a self-prompting module that generates multiple nearby viewpoints. With only several multi-view condition images, we can improve the geometric and visual consistency of Zero-1-to-3, which is cheap and effective.
  • Our proposed Cascade-Zero123 has better capability to handle complex scenes, even scenarios that were challenging for Zero-1-to-3, such as complex insects, humans, robots, or multiple objects stacked together. Our model has shown evident performance improvements on these previously difficult-to-solve scenes.
@@ 2.1 Single Image to 3D
Many researchers have studied the tasks of generating 3D models and achieving novel view synthesis using only a single image. Some researchers directly train a 3D model on 3D data [4, 63, 22, 38, 21], but they tend to have good generation quality only on scenes similar to the training set.
Recently, by constructing a conditional latent diffusion model based on camera viewpoints, many works made it possible to pre-train a single image-to-3D model [29, 51, 56, 64, 28, 34, 43, 15, 52, 27, 53, 50, 31, 32, 67, 61, 17, 66, 60]. Zero-1-to-3 [29] learns from large-scale multi-view images [10] to build the geometric priors of large-scale diffusion models. Zero-1-to-3 can lift various images that training sets have never been seen before to 3D with good quality. After that, many works have utilized Zero-1-to-3 as a module to enhance the quality of meshes or 3D models. Magic123 [43] combines the capabilities of Zero-1-to-3 and stable diffusion together to generate 3D models. One-2-3-45 [28] also leverages Zero-1-to-3 to generate different views to assist in mesh generation. However, these methods simply used it as a pre-trained 3D diffusion model tool with fixed checkpoints.
There have also been efforts to improve Zero-1-to-3. Approaches like Consistent1-to-3 [67], and SyncDreamer [31] aim to enhance the consistency of view generation by introducing priors during the denoising process in view-conditioned diffusion models. However, they add models like additional Transformers to render new views which will result in poor generalization due to these additional model’s capability, especially when facing unseen single images, so they may not integrate well with generalized latent diffusion models. In addition, the training of additional models usually costs a lot.
@@ 2.2 Multi-stage Diffusion Models
Cascade networks have been widely used in boosting the performance of models in the computer vision field, such as cascade RCNN [2], cascade DETR [68], and so on [25, 24]. After large-scale latent diffusion models became popular, many influential works also used cascade networks to improve the generation quality of models. For example, DeepFloyd-IF [47] achieved high-resolution and high-detail image generation by constructing three cascade pixel diffusion modules: a base model that generates a 64$\times$64 pixel image based on text prompts, and two super-resolution models. Furthermore, SDXL [41] used a cascade structure, in which a base model aligned with a refiner in the latent space to generate higher-quality and higher-resolution images. I2VGen-XL [70] also employed cascade diffusion models to enhance the quality of video generation. However, in the field of single image to 3D, there haven’t been works that utilize cascade structure to improve the quality of 3D generation. How to use cascade networks to build 3D consistency and improve 3D model generation quality has become an unsolved problem.
@@ 3.1 Preliminary
$q\left(x^{t}\mid x^{t-1}\right)=\mathcal{N}\left(x^{t};\sqrt{\alpha_{t}}x^{t-1},\left(1-\alpha_{t}\right)\mathbf{I}\right)$ (1)
where $\alpha$ is a scheduling hyper-parameter and $t\in[1,1000]$ denotes the diffusion timestep. $q\left(x^{t}\mid x^{t-1}\right)$ is the probability of estimating $x^{t}$ using $x^{t-1}$, and $\mathbf{I}$ is the normally distributed variance.
@@
Using a latent diffusion model with an encoder $f_{\phi}$, a denoiser U-Net $\epsilon_{\theta}$ and a decoder $\mathcal{D}$, the denoising process can be defined as follows.
@@
At the diffusion time step $t$, Zero-1-to-3 encodes the embedding of the input view and relative camera pose as $c\left(x_{c},\Delta R^{ct},\Delta T^{ct}\right)$ and set the objective loss function as
$L(\theta)=$ $\min_{\theta}\mathbb{E}_{z\sim\mathcal{E}(x),t,\epsilon\sim\mathcal{N}(0,1)}$ (4)
$\left\|\epsilon-\epsilon_{\theta}\left(z_{t},t,c\left(x_{c},\Delta R^{ct},\Delta T^{ct}\right)\right)\right\|_{2}^{2},$
@@ 3.2 Cascade-Zero123 Framework
Modeling the view rotation from a completely unseen image to a large rotation angle is highly challenging for Zero-1-to-3 [29]. We ease this task with a cascade structure, which progressively moves the camera pose with slight changes.
As shown in Fig. 2. Cascade-Zero123 consists of two cascade Zero-1-to-3 models. The first Zero-1-to-3 is referred to as Base-0123, and the second one is referred to as Refiner-0123. In terms of the structure, Base-0123 is responsible for generating multi-view images. Although these images lack consistency, they serve as rough condition inputs, which, along with their corresponding camera poses, are fed into the network of the Refiner-0123 model. The Refiner-0123 is aware of the multi-view input and it will compute the rotation and translation of different viewpoints to the final target view, which can enhance geometric and visual consistency in the outputs. After that, the training framework incorporates the model parameters of the second Refiner-0123 model back into the Base-0123 through exponential moving average (EMA).
Compared with Eq. 2, we can formulate the proposed Cascade-Zero123 as
$(\displaystyle x_{T}=f_{\phi_{r}}($$(x_{c}^{0},f_{\phi_{b}}(x,\Delta R_{b}^{c\{1:P\}},\Delta T_{b}^{c\{1:P\}})),$ (5)
$\Delta R_{r}^{\{1:P\}t},\Delta T_{r}^{\{1:P\}t}),$
where $f_{\phi_{b}}$ and $f_{\phi_{r}}$ are the Base-0123 and Refiner-0123 respectively, and the parameters of $f_{\phi_{b}}$ can be denoted as $\phi_{b}$ and those parameters of $f_{\phi_{r}}$ can be denoted as $\phi_{r}$. $x_{T}$ is the final target image. $P$ is the number of the prompt views, $\Delta R_{r}^{it},\Delta T_{r}^{it}$ is the prompted views camera pose of Refiner-0123.
@@ 3.3 Base-0123 Framework
In particular, at the beginning of the pre-training process, a set of input images denoted as $x_{c}^{0}$ is provided to Base-0123. Then we sample some rotation and translation matrices from the nearby viewpoints of the input view, which we denote as $\{(R^{i},T^{i})|i\in[1,P]\}$, where $P$ is the total number of prompt nearby views. We set the poses of $P$ viewpoints as constant ones to avoid the gap between the training and inference phases.
Our ablation experiments have also demonstrated that generating prompt images with small angles produces good multi-view consistency and final performance. Detailed view setting and hyper-parameter selection can be found in the implementation details of Sec. 4.2.
All the $P$ prompted viewpoints attended with the target view pose $(R_{t},T_{t})$ are concatenated with the same input embedding as $c\left(x_{c}^{0},\Delta R_{b}^{ci},\Delta T_{b}^{ci}\right)$, $i\in[1,P]$. Specifically, we calculate
$(\Delta R_{b}^{ci},\Delta T_{b}^{ci})=(R^{i},T^{i})\ominus(R^{c},T^{c})\;i\in[1,P]$ (6)
@@ 3.4 Refiner-0123 Framework
$(\Delta R_{r}^{it},\Delta T_{r}^{it})=(R^{t},T^{t})\ominus(R^{i},T^{i})\;i\in[1,P]$ (8)
@@
$(\hat{x}_{t}^{P},\Delta R_{r}^{Pt},\Delta T_{r}^{Pt}),(\hat{x}_{t}^{t},\Delta R_{r}^{tt},\Delta T_{r}^{tt}))$
@@
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V$ (10)
$Q=W_{Q}^{(i)}\cdot\varphi_{i}\left(z_{t}\right),K=W_{K}^{(i)}\cdot\tau_{\phi_{r}}(y),V=W_{V}^{(i)}\cdot\tau_{\phi_{r}}(y)$ (11)
@@
$\min_{\theta}$ $\mathbb{E}_{z\sim\mathcal{E}(x),t,\epsilon\sim\mathcal{N}(0,1)}\left\|\epsilon-\right.$ (12)
$\left.\epsilon_{\phi_{r}}\left(z_{r},t,c(x_{c}^{0},\hat{x}_{t}^{\{1:p\}},\Delta R_{r}^{\{1:p\}},\Delta T_{r}^{\{1:p\}})\right)\right\|_{2}^{2}.$

### v2

@@ Abstract
Synthesizing multi-view 3D from one single image is a significant but challenging task. Zero-1-to-3 methods have achieved great success by lifting a 2D latent diffusion model to the 3D scope. The target-view image is generated with a single-view source image and the camera pose as condition information. However, due to the high sparsity of the single input image, Zero-1-to-3 tends to produce geometry and appearance inconsistency across views, especially for complex objects. To tackle this issue, we propose to supply more condition information for the generation model but in a self-prompt way. A cascade framework is constructed with two Zero-1-to-3 models, named Cascade-Zero123, which progressively extract 3D information from the source image. Specifically, several nearby views are first generated by the first model and then fed into the second-stage model along with the source image as generation conditions. With amplified self-prompted condition images, our Cascade-Zero123 generates more consistent novel-view images than Zero-1-to-3. Experiment results demonstrate remarkable promotion, especially for various complex and challenging scenes, involving insects, humans, transparent objects, and stacked multiple objects etc. More demos and code are available at https://cascadezero123.github.io.
Keywords:
Single Image to 3D Novel View Synthesis Cascade Networks Diffusion Models
Figure 1: Rather than adopting limited input information, which Zero-1-to-3 [43] generation pipeline only has a single-view source image, Cascade-Zero123 progressively extracts the 3D information from more condition images by self-prompting. View-consistent images can be generated in a cascade manner. Cascade-Zero123 shows the strong capability on various complex objects, e.g. insects, robots, or multiple objects stacked.
@@ 1 Introduction
Generating the 3D object from a single image has become an appealing research topic, for its flexibility and convenience of creating 3D assets, which can be applied to various real-world applications including virtual reality, computer gaming, movie and animation production [56, 64, 113, 49, 82, 31, 76, 77, 13, 10, 97, 44, 38, 57]. However, due to the input information is too limited – only one image, it is quite challenging to reconstruct a high-fidelity 3D object. Many research works attempt to lift strong generation power from 2D large latent diffusion models [65, 53, 67, 63] to the 3D fields [39, 84, 8, 87, 98, 72, 112, 3, 29, 54, 21, 20, 48, 45, 92, 101, 105, 99, 19, 94, 58, 2, 91, 109], bringing possibilities to untractable 3D generation tasks.
As a representative work, Zero-1-to-3 [43] innovatively proposes to generate novel view images by conditioning the diffusion model with the source image and corresponding camera pose transition. The Zero-1-to-3 diffusion model is tuned on a large-scale multi-view 3D dataset, Objaverse [16, 15], and has achieved remarkable success and generalizability in synthesizing novel views from any single image. However, still suffering from the information sparsity of a single image, Zero-1-to-3 may inevitably fail to keep geometry or appearance consistency across different views, especially for objects with complex structures. As shown in Fig. 2, when the camera pose changes drastically, e.g. rotating more than $45^{\circ}$, it is hard to guarantee a high synthesis quality.
Providing more input information beyond a single image will undoubtedly ease the difficulty of generating 3D objects. We expect to improve the Zero-1-to-3 model by supplying more condition images from different views but in a self-prompting way. The main idea is to design a cascade framework to achieve this goal, which contains two Zero-1-to-3 models, i.e. Base-0123 and Refiner-0123. The Base-0123 model is conditioned to generate multiple images from various views. These images are constrained in a small range, e.g. $45^{\circ}$ in our implementation, to guarantee reliable quality as observed in Fig. 2. The Refiner-0123 model takes in all these generated images along with the source image to output the target-view image. We name the proposed framework as Cascade-Zero123. Rather than directly generating the target view as in Zero-1-to-3, Cascade-Zero123 lowers the synthesis difficulty by progressively digging out the 3D information starting from the input source image. To further enhance the Base-0123 model, we propose to update its parameters with a self-distillation manner from Refiner-0123. As shown in Fig. 2 and Fig. 1, our method can consistently promote the synthesis quality in all the transition ranges of camera poses.
@@
  • We propose a cascade framework to progressively extract 3D information from a single image, preventing inconsistency caused by drastic view transitions.
  • We design a self-prompting method that generates multiple nearby viewpoints, providing reliable condition information for the second-stage Refiner-0123 model.
  • Our proposed Cascade-Zero123 has shown notable promotion on complex scenes that are challenging for Zero-1-to-3, such as insects, humans, robots, or multiple objects stacked together.
Figure 2: The performance comparison of Zero-1-to-3 [43] and our methods on Google Scanned Object [18] with different camera pose rotation angles. When the camera pose changes drastically, the synthesis quality of Zero-1-to-3 will drop drastically. But our method can promote the synthesis quality in all the transition ranges of camera poses.
@@ 2.1 Single Image to 3D
Many researchers have studied the tasks of generating 3D models and achieving novel view synthesis using only a single image [90, 75, 110, 32, 37, 93, 41, 107, 12, 74] and text [68, 26, 51, 34, 52, 83, 14, 78, 60, 104, 108, 7]. Some researchers directly train a 3D model on 3D data [6, 95, 33, 54, 30, 62, 25], but they tend to have good generation quality only on scenes similar to the training set.
Recently, by constructing a conditional latent diffusion model based on camera viewpoints, many works have made it possible to pre-train a single image-to-3D model [43, 70, 79, 96, 42, 50, 61, 22, 71, 40, 73, 69, 46, 47, 102, 89, 24, 100, 88, 80]. Zero-1-to-3 [43] learns from large-scale multi-view images [16] to build the geometric priors of large-scale diffusion models. Zero-1-to-3 can lift various images that training sets have never been seen before to 3D with good quality. After that, many works have utilized Zero-1-to-3 as a module to enhance the quality of meshes or 3D models. Magic123 [61] combines the capabilities of Zero-1-to-3 and stable diffusion together to generate 3D models. One-2-3-45 [42] also leverages Zero-1-to-3 to generate different views to assist in mesh generation. However, these methods simply use it as pre-trained 3D diffusion model tools with fixed checkpoints.
There have also been efforts to improve Zero-1-to-3. Approaches like Consistent1-to-3 [102], aim to enhance the consistency of view generation by introducing priors during the denoising process in view-conditioned diffusion models. However, they add models like additional Transformers to render new views which will result in poor generalization due to these additional model’s capability, they may not integrate well with generalized latent diffusion models. In addition, the training of additional models usually costs a lot.
@@ 2.2 Multi-stage Diffusion Models
Cascade networks have been widely used in boosting the performance of models in the computer vision field, such as cascade RCNN [4], cascade DETR [103], and so on [36, 35]. After large-scale latent diffusion models become popular, many influential works also use cascade networks to improve the generation quality of models. For example, DeepFloyd-IF [66] achieves high-resolution and high-detail image generation by constructing three cascade pixel diffusion modules: a base model that generates a 64$\times$64 pixel image based on text prompts, and two super-resolution models. Furthermore, SDXL [59] uses a cascade structure, in which a base model aligned with a refiner in the latent space to generate higher-quality and higher-resolution images. I2VGen-XL [111] also employs cascade diffusion models to enhance the quality of video generation. Inspired by these cascade network designs, we attempt to progressively extract 3D information from a single image to improve the novel view synthesis quality.
@@ 3.1 Preliminary
$q\left(x^{t}\mid x^{t-1}\right)=\mathcal{N}\left(x^{t};\sqrt{\alpha_{t}}x^{t-1},\left(1-\alpha_{t}\right)\mathbf{I}\right),$ (1)
where $\alpha$ is a scheduling hyper-parameter and $t\in[1,1000]$ denotes the diffusion timestep. $q\left(x^{t}\mid x^{t-1}\right)$ estimate the probability of $x^{t}$ using $x^{t-1}$, and $\mathbf{I}$ is the normally distributed variance.
@@
Using a latent diffusion model with an encoder $f_{\phi}$, a denoiser U-Net $\epsilon_{\theta}$ and a decoder $\mathcal{D}$, the denoising process can be defined as follows:
@@
At the diffusion time step $t$, Zero-1-to-3 encodes the embedding of the input view and relative camera pose as $c\left(x_{c},\Delta R^{ct},\Delta T^{ct}\right)$ and set the objective loss function as:
$L(\theta)=\min_{\theta}\mathbb{E}_{z\sim\mathcal{E}(x),t,\epsilon\sim\mathcal{N}(0,1)}\left\|\epsilon-\epsilon_{\theta}\left(z_{t},t,c\left(x_{c},\Delta R^{ct},\Delta T^{ct}\right)\right)\right\|_{2}^{2},$ (4)
@@ 3.2 Cascade-Zero123 Framework
Suffering from the information sparsity of a single image, it is highly challenging for Zero-1-to-3 [43] to generate geometry and appearance consistent novel views with a single source image. We ease this task with a cascade structure, which progressively moves the camera pose with slight changes.
As shown in Fig. 3. Cascade-Zero123 consists of two cascade Zero-1-to-3 models. The first Zero-1-to-3 is referred to as Base-0123, and the second one is referred to as Refiner-0123. In terms of the structure, Base-0123 is responsible for generating multi-view images. These images serve as rough condition inputs, which, along with their corresponding camera poses, are fed into the network of the Refiner-0123 model. The Refiner-0123 is aware of the multi-view input and it will compute the rotation and translation of different viewpoints to the final target view. Supplying more condition images from different views to Refiner-0123 can ease the difficulty of generating novel views. After that, the training framework incorporates the model parameters of the second Refiner-0123 model back into the Base-0123 through exponential moving average (EMA).
Compared with Eq. 2, we can formulate the proposed Cascade-Zero123 as:
$(\displaystyle x_{T}=f_{\phi_{r}}($$(x_{c}^{0},f_{\phi_{b}}(x,\Delta R_{b}^{c\{1:P\}},\Delta T_{b}^{c\{1:P\}})),\Delta R_{r}^{\{1:P\}t},\Delta T_{r}^{\{1:P\}t}),$ (5)
where $f_{\phi_{b}}$ and $f_{\phi_{r}}$ are the Base-0123 and Refiner-0123 respectively, and the parameters of $f_{\phi_{b}}$ can be denoted as $\phi_{b}$ and those parameters of $f_{\phi_{r}}$ can be denoted as $\phi_{r}$. $x_{T}$ is the final target image. $P$ is the number of the prompt views, $\Delta R_{r}^{it},\Delta T_{r}^{it}$ is the prompted views camera pose of Refiner-0123. For easy understanding, we include Table 1 to list some symbols mentioned in the paper along with their meanings.
Table 1: Table of mentioned symbol representation of this paper.
Symbol Meaning
$x_{c}$, $x_{t}$ source image and target image
$\Delta R^{ct}$, $\Delta T^{ct}$ rotation and translation from the condition pose to the target pose
$\Delta R_{b},\Delta T_{b}$ rotation and translation condition to Base-0123
$\Delta R_{r},\Delta T_{r}$ rotation and translation condition to Refiner-0123
$f_{\phi_{b}}$, $f_{\phi_{r}}$ Base-0123 model and Refiner-0123 model
$\ominus$ pose transitions in the world coordinate
$\hat{x}_{t}^{1},...,\hat{x}_{t}^{P}$ self-prompted views
@@ 3.3 Base-0123 Framework
In particular, at the beginning of the pre-training process, a set of input images denoted as $x_{c}^{0}$ is provided to Base-0123. Then we sample some rotation and translation matrices from the nearby viewpoints of the input view, which we denote as $\{(R^{i},T^{i})|i\in[1,P]\}$, where $P$ is the total number of prompt nearby views. We set the poses of $P$ viewpoints as constant ones to avoid the gap between the training and inference phases. Constant poses can also save computation resources due to that these poses can be reused across different targets.
Our ablation experiments have also demonstrated that generating prompt images with small angles produces good performance. Detailed view setting and hyper-parameter selection can be found in the implementation details of Sec. 4.2.
All the $P$ prompted viewpoints attended with the target view pose $(R_{t},T_{t})$ are concatenated with the same input embedding as $c\left(x_{c}^{0},\Delta R_{b}^{ci},\Delta T_{b}^{ci}\right)$, $i\in[1,P]$. Specifically, we calculate:
$(\Delta R_{b}^{ci},\Delta T_{b}^{ci})=(R^{i},T^{i})\ominus(R^{c},T^{c})\;i\in[1,P],$ (6)
@@ 3.4 Refiner-0123 Framework
$(\Delta R_{r}^{it},\Delta T_{r}^{it})=(R^{t},T^{t})\ominus(R^{i},T^{i})\;i\in[1,P],$ (8)
@@
$(\hat{x}_{t}^{P},\Delta R_{r}^{Pt},\Delta T_{r}^{Pt}),(\hat{x}_{t}^{t},\Delta R_{r}^{tt},\Delta T_{r}^{tt})).$
@@
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V.$ (10)
$Q=W_{Q}^{(i)}\cdot\varphi_{i}\left(z_{t}\right),K=W_{K}^{(i)}\cdot\tau_{\phi_{r}}(y),V=W_{V}^{(i)}\cdot\tau_{\phi_{r}}(y).$ (11)
@@
$\min_{\theta}\mathbb{E}_{z\sim\mathcal{E}(x),t,\epsilon\sim\mathcal{N}(0,1)}\left\|\epsilon-\right.\left.\epsilon_{\phi_{r}}\left(z_{r},t,c(x_{c}^{0},\hat{x}_{t}^{\{1:p\}},\Delta R_{r}^{\{1:p\}},\Delta T_{r}^{\{1:p\}})\right)\right\|_{2}^{2}.$ (12)